Grade 11Mechanics

y=25cos (3.14t+40x)....from this plane progressive wave determine the time period and wavelength

Prithvi singh
9 Years agoGrade 11

1 Answer

Khimraj
9 Years ago
In equation y = Acos(wt+ kx) A = amplitude of particlew = 2π/Tk =2π/λSo time period = 2π/3.14 = 2 secWavelength = 2π/40 cm